# Observations of high-frequency spectral peaks from in-situ waves in ice data: evidence for nonlinear waves in ice triad interactions?

### v1

@@
Observations of high-frequency spectral peaks from in-situ waves in ice data: evidence for nonlinear waves in ice triad interactions?
Jean Rabault
Norwegian Meteorological Institute
Oslo, Norway
jean.rblt@proton.me
&Joey Voermans
The University of Melbourne
Melbourne, Australia
jvoermans@unimelb.edu.au
&Takehiko Nose
The University of Tokyo
Kashiwa, Japan
tak.nose@edu.k.u-tokyo.ac.jp
&Graig Sutherland
Environment Climate Change Canada
Dorval, Canada
graig.sutherland@gmail.com
&Alexander Babanin
The University of Melbourne
Melbourne, Australia
a.babanin@unimelb.edu.au
&Takuji Waseda
The University of Tokyo
Kashiwa, Japan
waseda@edu.k.u-tokyo.ac.jp
&Tsubasa Kodaira
The University of Tokyo
Kashiwa, Japan
kodaira@edu.k.u-tokyo.ac.jp
&Atle Jensen
University of Oslo
Oslo, Norway
atlej@math.uio.no
&Lars Willas Dreyer
University of Oslo
Oslo, Norway
larswd@math.uio.no
&Øyvind Breivik
Norwegian Meteorological Institute
Bergen, Norway
oyvindb@met.no
&Gaute Hope
Norwegian Meteorological Institute
Bergen, Norway
gauteh@met.no
&Malte Müller
Norwegian Meteorological Institute
Oslo, Norway
maltem@met.no
&Zhaohui Cheng
Uppsala University
Uppsala, Sweden
zhaohui.cheng@geo.uu.se
&Lichuan Wu
Uppsala University
Uppsala, Sweden
lichuan.wu@geo.uu.se
&Aleksey Marchenko
The University Centre in Svalbard
Svalbard, Spitsbergen, Norway
alekseym@unis.no
&Petra Heil
University of Tasmania
Hobart, Australia
petra.heil@utas.edu.au &Karsten Trulsen
University of Oslo
Oslo, Norway
karstent@math.uio.no
@@ 1 Introduction
The Marginal Ice Zone (MIZ) is the area of sea ice that is strongly influenced by the propagation of open ocean waves. These waves can, in particular, lead to breakup of sea ice collins2015situ ; voermans2020experimental , which can dramatically alter sea ice conditions and greatly influences sea ice drift sutherland2022estimating ; thomson2022wave and melting thomson2014swell ; roach2019advances .
@@
Numerous papers have discussed the dispersion relation for waves in ice; see, e.g. squire1995ocean ; voermans2021wave . Often, the real (propagation properties) and imaginary (attenuation) parts of the dispersion relation are considered separately from each other, in which case $k=|\mathbb{k}_{r}|$ stands for the real part of the wavenumber, and one may write $\mathbb{k}$ instead of $\mathbb{k}_{r}$ as a shorthand notation. For example, greenhill1886wave wrote down (though it may have been known even earlier) a now classical form for the real part of the wave dispersion relation in ice under the assumption that the ice acts as a thin elastic sheet where flexural stiffness effects are present:
@@
Interestingly, the term representing the effect of elasticity (i.e. $Bk^{5}$) makes it easy for the nonlinear resonant wave triad necessary conditions between 3 wave components (1,2,3)(1,2,3)( 1 , 2 , 3 ), i.e. (using the shorthand notation mentioned above and omitting the r indexes for simplicity):
@@
to be fulfilled phillips1966dynamics ; phillips1981wave ; madsen1993bound , as we will discuss later in this manuscript. Although triads and other resonant interactions have, to the best of our knowledge, not been discussed extensively on the basis of field measurements, a few theoretical works can be found that discuss their possible existence marchenko1999stability ; marchenko1999stability2 . The only reported observations of high-frequency spectral peaks we know of, and that may be similar to ours, are the ones from johnson2021observing , though these seem to match a 4th order, rather than 2nd order (as we observe here, see below) high-frequency spectral peak, and it is uncertain what their origin is and whether they may come from the effect of wind per se, or can be the result of some nonlinear wave energy redistribution.
Regarding wave attenuation, the most widely accepted consensus, based on theoretical arguments, laboratory experiments, and empirical measurements in the ice wadhams1973attenuation ; yu2022 ; squire2020ocean ; loken2021wave ; zhao2015modeling ; zhao2015wave , is that the amount of wave energy dissipation at each frequency is proportional to the local amount of wave energy. This leads to an exponential attenuation of the waves as they propagate in the sea ice and is consistent with the wavenumber formalism introduced above, resulting in:
@@
Generally, $k_{i}(f)$ is an increasing function of wave frequency $f$ meylan2014situ ; yu2022 . Currently, there is general agreement in the literature on the existence of a scaling $k_{i}(f)\propto f^{n}$, with the value of the exponent $n$ reported to be typically around 4.5 yu2022 , although there is some variability in the literature depending on the exact choices for the wave in ice damping parameterization and the dataset(s) over which this relation is fitted.
@@
However, there is still a debate in the literature on what physical mechanism, or combination of mechanisms, physically determines the attenuation coefficient $k_{i}(f)$, with effects such as scattering bennetts2010three ; kohout2008elastic ; montiel2016attenuation ; zhao2016diffusion ; bennetts2024thin , viscous dissipation zhao2015modeling ; sutherland2019two ; rabault2017measurements ; marchenko2018wave , turbulence voermans2019wave ; smith2020pancake , floe-floe interaction herman2019wave ; loken2022experiments ; herman2019wave ; smith2020pancake , viscoelasticity in ice mosig2015comparison ; zhang2021theoretical ; zhao2018three ; marchenko2021laboratory , and overwash pitt2022model ; zha2023numerical , regularly discussed. Although several formulations have been proposed and implemented to parameterize waves in ice damping (for example, at least 10 different parameterizations are available in WaveWatch III, tolman1991third ; rogers2013implementation ; KGulla ), and progress has been reported recently using fitted parameterizations based on scaling laws to unify data sets that span several orders of magnitude of waves in ice parameters yu2022 , models are still largely imperfect at representing waves in ice and, in particular, the wave in ice attenuation $k_{i}(f)$ KGulla ; rabault2024situ . Possibly, different combinations of dissipation mechanisms may play a role in different wave and ice conditions, making the problem complex and difficult to disentangle and solve in a general way herman2024apparent . If triads exist for waves in ice, this may add an additional possible dissipation mechanism by stealing energy from the main lower-frequency spectral components that are weakly attenuated and moving it into higher frequency components where energy dissipation is stronger.
@@
The ongoing discussions and uncertainties about the details of the mechanisms causing the wave attenuation and the exact formula to use highlight the need for more ground-truth direct in-situ observations to help constrain and calibrate theoretical models. This has been the focus of the waves in ice community for many decades wadhams1973attenuation ; doble2006wave ; thomson2013sea ; kohout2016situ , and our groups have recently contributed significant amounts of such data rabault2023dataset ; rabault2024position ; muller2025distributed ; rabault2022tracking using new generations of open source, low-cost instruments based on the OpenMetBuoy (OMB) and closely related line of designs rabault2016measurements ; rabault2017measurements ; rabault2020open ; marchenko2019wave ; loken2021wave ; rabault2022openmetbuoy ; kodaira2024affordable ; kodaira2022development ; hope2025sfy ; MoreRoomattheTopHowSmallBuoysHelpRevealtheDetailedDynamicsoftheAirSeaInterface . The availability of larger in-situ datasets allows, in turn, to perform more detailed and representative analysis of real-world ground-truth data, which has contributed to reveal interesting dynamics dreyer2024direct ; rabault2024situ ; nose2024observation ; nose2023comparison .
@@ 2.1.1 Timeseries data: Tempelfjorden 2015 (T15)
The first and most comprehensive dataset comes from the 2015 waves in ice measurements performed on landfast ice in Tempelfjorden, Svalbard, using a MOXA logger sutherland2016observations ; rabault2016measurements . A total of 5 Vectornav VN100 IMU (Inertial Motion Unit)-based sensors were deployed. Of these, the data from three locations have proved to work well from a technical point of view and produce useful data without measurement dropouts: these correspond to the IMUs with IDs VN1005, VN1002, and VN1007 (from closest to the open water to deepest into the ice), which we will consider in the following.
The VN100 data acquired during this deployment contain the raw X, Y and Z (that is, in the reference frame of the VN100 IMUs) sea ice motion data. In addition, the data are obtained by point-wise sampling of the internal readings of individual sensors at 10 Hz, while the sensors themselves run internally at 800 Hz. This means that we need to apply Kalman filtering a posteriori to the data sutherland2017method (since the sensors frame of reference is used) and that the noise level is slightly higher than in other deployments (since sampling rather than averaging is used). The data and processing are openly available; see the details in Appendix A.
@@ 2.1.2 Timeseries data: Yermak plateau 2020 (Y20)
The second dataset comes from the OMB-v2018 deployment performed in 2020 over the Yermak plateau from the hovercraft R/H Sabvabaa rabault2023dataset . There, a total of six OMB-v2018s, each containing a VN100 IMU, were deployed on sea ice. Later, one of the buoys, with ID 18954, was found and recovered after being stranded on the northwest coast of Iceland dreyer2024direct . Therefore, the raw timeseries of the VN100 motion data recorded by this instrument are available from the on-board SD card. The IMU configuration is improved compared to the T15 dataset. More specifically, these VN100 IMU data contain Kalman-filtered (by the processor on board of the IMU package) 10 Hz averaged (by performing an 80-point averaging, instead of individual sample extraction, which provides lower noise) acceleration data in the Earth frame of reference rabault2020open ; dreyer2024direct . The data are openly available; see the details in Appendix A.
@@ 2.1.3 Spectra over Iridium
The third dataset includes data from Sofar Spotter raghukumar2019directional , OMB-v2018 rabault2020open , and OMB-v2021 rabault2022openmetbuoy deployments, which collected data in both the Arctic and Antarctic. In these deployments, the buoys could not be recovered, so only the 1-dimensional power spectral density (PSD) transmitted over iridium satellite communications is available.
@@
  • The Davis 2020 (D20) Antarctic deployment rabault2023dataset . In this deployment, two OMB-v2018 and two Sofar Spotter were deployed on Antarctic landfast ice along a transect perpendicular to the broken ice edge, close to the Davis Research Station.
@@
  • The Yermak 2020 Arctic deployment. This is the same deployment as Y20 discussed above rabault2023dataset , but data from other buoys are considered.
  • The 2022 CIRFA-UIT Arctic deployment. This deployment was performed in the Western Fram straight, West of Greenland. 19 OMB-v2021s were deployed from the R/V Kronprins Haakon. The instruments were deployed in April and May 2022, and data transmissions took place until December 2022 rabault2024position .
  • The 2022 AWI-UTOKYO Arctic deployment. This deployment was performed with 15 OMB-v2021s. The OMBs were deployed on the sea ice northwest of Svalbard in July 2022 and transmitted until October 2022 rabault2024position .
@@ 2.2.1 Kalman processing of the T15 data
For the T15 data, Kalman processing must be applied to the raw IMU data measured in the IMU reference frame to obtain the ice motion in the Earth (North, East, Down) reference frame. This is done using the same open source Kalman filter that is used on the OMB-v2021 rabault2022openmetbuoy . The Kalman filter itself is available at https://github.com/adafruit/Adafruit_AHRS and https://github.com/gauteh/ahrs-fusion (both accessed 2025-07), and all code and instructions are available along with the data released with this manuscript; see Appendix A.
@@ 2.2.2 Power Spectral Density calculation
The wave Power Spectral Density (PSD) is calculated from the downward acceleration component in the Earth reference frame. When computing the PSD, we use the Welch method welch1967use , typically with an FFT length of 2048 points, 50% overlap, and a Hann windowing function. This is a standard method for generating the PSD, and is typically similar to what is used in, e.g., the firmware running on board of the OMB-v2018 and OMB-v2021 rabault2020open ; rabault2022openmetbuoy , and similar buoy designs kohout2015device . We also performed consistency checks slightly modifying these parameters during the exploration phase of this work, which confirmed that the results were not affected.
High-frequency peaks in the PSD of the wave motion are easier to visualize in the acceleration spectra rather than the elevation spectra. This is because, at the hardware level, the actual physical sensors used in the IMUs measure the acceleration, not the elevation, of the waves. Therefore, converting from the acceleration spectrum to the elevation spectrum requires applying the scaling $\mathrm{PSD}_{\eta}=\omega^{-4}\mathrm{PSD}_{\mathrm{acc}}$, where $\mathrm{PSD}_{\eta}$ is the PSD of the elevation and $\mathrm{PSD}_{\mathrm{acc}}$ is the PSD of the acceleration tucker2001waves . Thus, the spectral noise level goes from a constant (horizontal line) in $\mathrm{PSD}_{\mathrm{acc}}$ to a sloped line in $\mathrm{PSD}_{\eta}$, which makes it harder to visually distinguish noise from signal in the elevation spectrum. As a consequence, we consistently show acceleration spectra rather than elevation spectra in the following.
@@ 2.2.3 Cross Spectral Density and dispersion relation calculation
The cross spectral density (CSD) between two adjacent sensors can be used to recover the frequency-dependent phase shift due to the wave propagation delay between the two sensors and hence the dispersion relation longuet1963observation ; tucker2001waves ; marchenko2017field . This analysis can be performed on the T15 dataset; for other datasets, we do not have data from 2 or more adjacent sensors.
When performing a CSD analysis between two sensors (here, generally indicated by 111 and 222) to obtain the dispersion relation, the following procedure is used:
  • The cross-spectral coherence is calculated following the formula: $\mathrm{CSC}_{1,2}(f)=\frac{|P_{1,2}(f)|^{2}}{P_{1,1}(f)P_{2,2}(f)}$, where $P_{i,j}(f)=\mathrm{FFT}_{i}(f)\mathrm{FFT}_{j}(f)^{*}$ is the cross-spectral density between the IMUs number $i$ and $j$, $\mathrm{FFT}$ indicates the Fast Fourier Transform operator, and indicates complex conjugate. In practice, this is evaluated using the Welch method, i.e. averaging over sub-segments, similar to what is done for the PSD. A coherence value significantly above 0 indicates that phase locking is observed between the two sensors for the frequency considered, and that it is meaningful to extract a phase shift value from the cross-spectral density $P_{1,2}$. Here, we fix the threshold for the cross-spectral coherence at a value of 0.25. This is somewhat low, but this choice is dictated by the low signal-to-noise ratio in the acceleration data, as visible in the PSDs, so that the signal-to-noise ratio of the CSC is also relatively limited. In the frequency domain around the high-frequency peak, where the signal-to-noise ratio goes down to around 5, we even consider a lower coherence value threshold of 0.15, though we acknowledge that this is pushing the limit and the results should be considered with care and only viewed as an indication. However, as visible in Fig. 3, the results are consistent and correspond well to the theory, giving us a posteriori confidence in these relatively low threshold values.

### v5

@@ 1 Introduction
The Marginal Ice Zone (MIZ) is the area of sea ice that is strongly influenced by the propagation of open ocean waves. These waves can, in particular, lead to breakup of sea ice (Collins III et al., 2015; Voermans et al., 2020), which can dramatically alter sea ice conditions and greatly influences sea ice drift (Sutherland et al., 2022; Thomson, 2022) and melting (Thomson and Rogers, 2014; Roach et al., 2019).
@@
Numerous papers have discussed the dispersion relation for waves in ice; see, e.g. Squire et al. (1995); Voermans et al. (2021). Often, the real (propagation properties) and imaginary (attenuation) parts of the dispersion relation are considered separately from each other, in which case $k=|\mathbb{k}_{r}|$ stands for the real part of the wavenumber, and one may write $\mathbb{k}$ instead of $\mathbb{k}_{r}$ as a shorthand notation. For example, Greenhill (1886) wrote down (though it may have been known even earlier) a now classical form for the real part of the wave dispersion relation in ice under the assumption that the ice acts as a thin elastic sheet where flexural stiffness effects are present:
@@
Interestingly, the term representing the effect of elasticity (i.e. $Bk^{5}$) makes it easy for the nonlinear resonant wave triad necessary conditions between 3 wave components $(1,2,3)$, i.e. (using the shorthand notation mentioned above and omitting the r indexes for simplicity):
@@
to be fulfilled (Phillips, 1966, 1981; Madsen and Sørensen, 1993), as we will discuss later in this manuscript. Although triads and other resonant interactions have, to the best of our knowledge, not been discussed extensively on the basis of field measurements, a few theoretical works can be found that discuss their possible existence (Marchenko, 1999b, ), and this was discussed in detail more recently by Pierce et al. (2024, 2025). In particular, Pierce et al. (2024) showed through both analytical and numerical simulations that triads can develop over a few tens of wavelengths in ideal conditions and extract significant amounts of energy from the main incoming wave. The only reported observations of high-frequency spectral peaks we know of, and that may be similar to ours, are the ones from Johnson et al. (2021), though these seem to match a 4th order, rather than 2nd order (as we observe here, see below) high-frequency spectral peak, and it is uncertain what their origin is and whether they may come from the effect of wind per se, or can be the result of some nonlinear wave energy redistribution.
Regarding wave attenuation, the most widely accepted consensus, based on theoretical arguments, laboratory experiments, and empirical measurements in the ice (Wadhams, 1973; Yu et al., 2022; Squire, 2020; Løken et al., 2021; Zhao et al., 2015; Zhao and Shen, 2015), is that the amount of wave energy dissipation at each frequency is proportional to the local amount of wave energy. This leads to an exponential attenuation of the waves as they propagate in the sea ice and is consistent with the wavenumber formalism introduced above, resulting in:
@@
Generally, $k_{i}(f)$ is an increasing function of wave frequency $f$ (Meylan et al., 2014; Yu et al., 2022). Currently, there is general agreement in the literature on the existence of a scaling $k_{i}(f)\propto f^{n}$, with the value of the exponent $n$ reported to be typically around 4.5 (Yu et al., 2022), although there is some variability in the literature depending on the exact choices for the wave in ice damping parameterization and the dataset(s) over which this relation is fitted.
@@
However, there is still a debate in the literature on what physical mechanism, or combination of mechanisms, physically determines the attenuation coefficient $k_{i}(f)$, with effects such as scattering (Bennetts et al., 2010; Kohout and Meylan, 2008; Montiel et al., 2016; Zhao and Shen, 2016; Bennetts et al., 2024), viscous dissipation (Zhao et al., 2015; Sutherland et al., 2019; Rabault et al., 2017; Marchenko, 2018), turbulence (Voermans et al., 2019; Smith and Thomson, 2020), floe-floe interaction (Herman et al., 2019; Løken et al., 2022; Herman et al., 2019; Smith and Thomson, 2020), viscoelasticity in ice (Mosig et al., 2015; Zhang and Zhao, 2021; Zhao and Shen, 2018; Marchenko et al., 2021), and overwash (Pitt et al., 2022; Zha et al., 2023), regularly discussed. Although several formulations have been proposed and implemented to parameterize waves in ice damping (for example, at least 10 different parameterizations are available in WaveWatch III, (Tolman, 1991; Rogers and Orzech, 2013; Gulla, 2025)), and progress has been reported recently using fitted parameterizations based on scaling laws to unify datasets that span several orders of magnitude of waves in ice parameters (Yu et al., 2022), models are still largely imperfect at representing waves in ice and, in particular, the wave in ice attenuation $k_{i}(f)$ (Gulla, 2025; Rabault et al., 2024a). Possibly, different combinations of dissipation mechanisms may play a role in different wave and ice conditions, making the problem complex and difficult to disentangle and solve in a general way (Herman, 2024). If triads exist for waves in ice, this may add an additional possible dissipation mechanism by stealing energy from the main lower-frequency spectral components that are weakly attenuated and moving it into higher frequency components where energy dissipation is stronger.
@@
The ongoing discussions and uncertainties about the details of the mechanisms causing the wave attenuation and the exact formula to use highlight the need for more ground-truth direct in-situ observations to help constrain and calibrate theoretical models. This has been the focus of the waves in ice community for many decades (Wadhams, 1973; Doble et al., 2006; Thomson et al., 2013; Kohout et al., 2016), and our groups have recently contributed significant amounts of such data (Rabault et al., 2023, 2024b; Müller et al., 2025a; Rabault, 2022) using new generations of open source, low-cost instruments based on the OpenMetBuoy (OMB) and closely related line of designs (Rabault et al., 2016b, 2017, 2020; Marchenko et al., 2019; Løken et al., 2021; Rabault et al., 2022; Kodaira et al., 2024, 2022; Hope et al., 2025; Cavaleri et al., 2025). The availability of larger in-situ datasets allows, in turn, to perform more detailed and representative analysis of real-world ground-truth data, which has contributed to reveal interesting dynamics (Dreyer et al., 2025; Rabault et al., 2024a; Nose et al., 2024, 2023).
@@ 2.1.1 Timeseries data: Tempelfjorden 2015 (T15)
The first and most comprehensive dataset comes from the 2015 waves in ice measurements performed on landfast ice in Tempelfjorden, Svalbard, using a MOXA logger (Sutherland and Rabault, 2016; Rabault et al., 2016b). A total of 5 Vectornav VN100 IMU (Inertial Motion Unit)-based sensors were deployed. Of these, the data from three locations have proved to work well from a technical point of view and produce useful data without measurement dropouts: these correspond to the IMUs with IDs VN1005, VN1002, and VN1007 (from closest to the open water to deepest into the ice), which we will consider in the following.
The VN100 data acquired during this deployment contain the raw X, Y and Z (that is, in the reference frame of the VN100 IMUs) sea ice motion data. In addition, the data are obtained by point-wise sampling of the internal readings of individual sensors at 10 Hz, while the sensors themselves run internally at 800 Hz. This means that we need to apply Kalman filtering a posteriori to the data (Sutherland et al., 2017b) (since the sensors frame of reference is used) and that the noise level is slightly higher than in other deployments (since sampling rather than averaging is used). The data and processing are openly available; see the details in Appendix A.
@@ 2.1.2 Timeseries data: Yermak plateau 2020 (Y20)
The second dataset comes from the OMB-v2018 deployment performed in 2020 over the Yermak plateau from the hovercraft R/H Sabvabaa (Rabault et al., 2023). There, a total of six OMB-v2018s, each containing a VN100 IMU, were deployed on sea ice. Later, one of the buoys, with ID 18954, was found and recovered after being stranded on the northwest coast of Iceland (Dreyer et al., 2025). Therefore, the raw timeseries of the VN100 motion data recorded by this instrument are available from the on-board SD card. The IMU configuration is improved compared to the T15 dataset. More specifically, these VN100 IMU data contain Kalman-filtered (by the processor on board of the IMU package) 10 Hz averaged (by performing an 80-point averaging, instead of individual sample extraction, which provides lower noise) acceleration data in the Earth frame of reference (Rabault et al., 2020; Dreyer et al., 2025). The data are openly available; see the details in Appendix A.
@@ 2.1.3 Spectra over Iridium
The third dataset includes data from Sofar Spotter (Raghukumar et al., 2019), OMB-v2018 (Rabault et al., 2020), and OMB-v2021 (Rabault et al., 2022) deployments, which collected data in both the Arctic and Antarctic. In these deployments, the buoys could not be recovered, so only the 1-dimensional power spectral density (PSD) transmitted over iridium satellite communications is available.
@@
  • The Davis 2020 (D20) Antarctic deployment (Rabault et al., 2023). In this deployment, two OMB-v2018 and two Sofar Spotter were deployed on Antarctic landfast ice along a transect perpendicular to the broken ice edge, close to the Davis Research Station.
@@
  • The Yermak 2020 Arctic deployment. This is the same deployment as Y20 discussed above (Rabault et al., 2023), but data from other buoys are considered.
  • The 2022 CIRFA-UIT Arctic deployment. This deployment was performed in the Western Fram straight, West of Greenland. 19 OMB-v2021s were deployed from the R/V Kronprins Haakon. The instruments were deployed in April and May 2022, and data transmissions took place until December 2022 (Rabault et al., 2024b).
  • The 2022 AWI-UTOKYO Arctic deployment. This deployment was performed with 15 OMB-v2021s. The OMBs were deployed on the sea ice northwest of Svalbard in July 2022 and transmitted until October 2022 (Rabault et al., 2024b).
@@ 2.2.1 Kalman processing of the T15 data
For the T15 data, Kalman processing must be applied to the raw IMU data measured in the IMU reference frame to obtain the ice motion in the Earth (North, East, Down) reference frame. This is done using the same open source Kalman filter that is used on the OMB-v2021 (Rabault et al., 2022). The Kalman filter itself is available at https://github.com/adafruit/Adafruit_AHRS and https://github.com/gauteh/ahrs-fusion (both accessed 2025-07), and all code and instructions are available along with the data released with this manuscript; see Appendix A.
@@ 2.2.2 Power Spectral Density calculation
The wave Power Spectral Density (PSD) is calculated from the downward acceleration component in the Earth reference frame. When computing the PSD, we use the Welch method (Welch, 1967), typically with an FFT length of 2048 points, 50% overlap, and a Hann windowing function. This is a standard method for generating the PSD, and is typically similar to what is used in, e.g., the firmware running on board of the OMB-v2018 and OMB-v2021 (Rabault et al., 2020, 2022), and similar buoy designs (Kohout et al., 2015). We also performed consistency checks slightly modifying these parameters during the exploration phase of this work, which confirmed that the results were not affected.
High-frequency peaks in the PSD of the wave motion are easier to visualize in the acceleration spectra rather than the elevation spectra. This is because, at the hardware level, the actual physical sensors used in the IMUs measure the acceleration, not the elevation, of the waves. Therefore, converting from the acceleration spectrum to the elevation spectrum requires applying the scaling $\mathrm{PSD}_{\eta}=\omega^{-4}\mathrm{PSD}_{\mathrm{acc}}$, where $\mathrm{PSD}_{\eta}$ is the PSD of the elevation and $\mathrm{PSD}_{\mathrm{acc}}$ is the PSD of the acceleration (Tucker and Pitt, 2001). Thus, the spectral noise level goes from a constant (horizontal line) in $\mathrm{PSD}_{\mathrm{acc}}$ to a sloped line in $\mathrm{PSD}_{\eta}$, which makes it harder to visually distinguish noise from signal in the elevation spectrum. As a consequence, we consistently show acceleration spectra rather than elevation spectra in the following.
@@ 2.2.3 Cross Spectral Density and dispersion relation calculation
The cross spectral density (CSD) between two adjacent sensors can be used to recover the frequency-dependent phase shift due to the wave propagation delay between the two sensors and hence the dispersion relation (Longuet-Higgins, 1963; Tucker and Pitt, 2001; Marchenko et al., 2017). This analysis can be performed on the T15 dataset; for other datasets, we do not have data from 2 or more adjacent sensors.
When performing a CSD analysis between two sensors (here, generally indicated by $1$ and $2$) to obtain the dispersion relation, the following procedure is used:
  • The cross-spectral coherence is calculated following the formula: $\mathrm{CSC}_{1,2}(f)=\frac{|P_{1,2}(f)|^{2}}{P_{1,1}(f)P_{2,2}(f)}$, where
    $P_{i,j}(f)=\mathrm{FFT}_{i}(f)\mathrm{FFT}_{j}(f)^{*}$ is the cross-spectral density between the IMUs number $i$ and $j$, $\mathrm{FFT}$ indicates the Fast Fourier Transform operator, and indicates complex conjugate. In practice, this is evaluated using the Welch method, i.e. averaging over sub-segments, similar to what is done for the PSD. A coherence value significantly above 0 indicates that phase locking is observed between the two sensors for the frequency considered, and that it is meaningful to extract a phase shift value from the cross-spectral density $P_{1,2}$. Here, we fix the threshold for the cross-spectral coherence at a value of 0.25. This is somewhat low, but this choice is dictated by the low signal-to-noise ratio in the acceleration data, as visible in the PSDs, so that the signal-to-noise ratio of the CSC is also relatively limited. In the frequency domain around the high-frequency peak, where the signal-to-noise ratio goes down to around 5, we even consider a lower coherence value threshold of 0.15, though we acknowledge that this is pushing the limit and the results should be considered with care and only viewed as an indication. However, as visible in Fig. 3, the results are consistent and correspond well to the theory, giving us a posteriori confidence in these relatively low threshold values.